# Edge Learning for 6G-enabled Internet of Things: A Comprehensive Survey of Vulnerabilities, Datasets, and Defenses

### v1

@@ II-D Quality assessment
In our study, we conducted a quality assessment using a questionnaire that included ten questions related to the research questions and objectives of the study. These questions covered a range of topics, including 1) the clarity and appropriateness of the research questions, 2) the comprehensiveness of the literature review, 3) the appropriateness and relevance of the datasets used, 4) the effectiveness of the classification methods used, 5) the depth and comprehensiveness of the discussions on proposed solutions, 6) the degree to which the proposed solutions are supported by experimental analysis, and 7) the clarity of the presentation of the results. By answering these questions, we were able to evaluate the quality and identify the advantages and disadvantages of each work. The formulation of the quality assessment criteria in Table III is based on eight questions that assess the quality of our research in relation to the research questions.
III An Overview of ML-Associated and 6G-enabled IoT Security
In the context of ML-Associated and 6G-enabled IoT Security, a threat refers to potential danger or harm to the security of the system, while a vulnerability refers to a weakness or flaw in the system that can be exploited by a threat actor [37]. Threats in ML-associated 6G-enabled IoT security inherit the threats associated with the previous generation communication networks including but not limited to data breaches, denial of service (DoS) attacks, unauthorized access to sensitive data, and tampering ML models [38]. On the other hand, vulnerabilities in ML-Associated and 6G-enabled IoT Security include outdated software, unsecured network connections, weak passwords, and insecure authentication mechanisms. These vulnerabilities can make it easier for a threat actor to carry out an attack on the system [39].
In this section, we present state-of-the-art reviews related to ML-related and 6G-enabled IoT security. We classify them into two categories: 1) ML-based security for 6G-IoT systems, which incorporate studies on IoT security and the vision of future 6G networks, and 2) ML-associated threats, which provide an overview of works on threats facing ML-associated paradigms, namely centralized, FL and distributed learning. In Table I, we provide a detailed comparison between our work and state-of-the-art studies.
III-A Machine Learning For 6G–IoT security
III-A1 Usage of ML Paradigms in 6G-enabled IoT Applications
The table IV provides an overview of how the different machine learning paradigms can be used in 6G-enabled IoT applications. The different machine learning paradigms can be categorized into five paradigms, including, Supervised Learning, Unsupervised Learning, Reinforcement Learning, Semi-Supervised Learning, and Transfer Learning [39, 40, 41]. Supervised learning is the most commonly used paradigm, where the model is trained on labeled data and learns to map inputs to outputs based on example pairs of input-output data. On the other hand, unsupervised learning involves training the model on unlabeled data and finding patterns and relationships in the data without explicit guidance. Semi-supervised learning combines both supervised and unsupervised learning, and the model is trained on both labeled and unlabeled data to improve its performance. Reinforcement learning involves training the model to make decisions based on feedback from the environment and taking actions that maximize reward over time.
III-A2 IoT-related security
Hussain et al. [12] explored the ways in which ML and Deep Learning (DL) have impacted the loT ecosystem from a security and privacy perspective. The survey first presents the background on the security and privacy concerns and challenges facing the IoT, including those obstacles related to the resource constraints associated with IoT devices, as well as the attack vectors and security expectations. This is followed by highlighting various ML and DL mechanisms and their applicability to IoT security in various scopes of applications such as forensic face recognition, cryptographic security character identification, and malicious code detection. In addition, the survey highlights the shortcomings that may be confronted by the adoption of ML techniques in IoT, including resource limitations, where IoT devices may not be suitable to support or execute sophisticated computational processing. Among other recommendations, the authors advocate that in order to overcome some of the limitations of ML approaches to IoT security, both DL and Deep Reinforcement Learning (DRL) theoretical frameworks should be further enhanced to allow adequate quantification of performances based on metrics such as computing complexity.
In the same context, Zaman et al. [18] provide a study regarding IoT security threats by layer, as well as security schemes to address them. The layers involved are perception, network, transport, processing, and application, with a focus on different IoT protocols related to each layer of the ecosystem. In response to the layer-based threats, rule-based (such as fuzzy logic) and AI-based layer response actions were presented, including ML (such as SVM) and DL (such as DNN and KNN) based systems, as well as performance evaluations of these AI-based layer-wise response actions.
In addition, Mohanta et al. [13] suggest the use of emerging technologies such as AI, ML, and blockchain technologies to address existing security and privacy issues in IoT applications, such as jamming, DoS, and malicious nodes identification. The authors begin by highlighting layer-wise security issues in the IoT system and then answer the question of "how can these technologies be used to mitigate security threats in IoT?" in an overview. A study by Sarker et al. [25] offered a holistic view of IoT security intelligence, which is driven by ML and DL techniques that mine information out of raw data to smartly safeguard IoT systems against a range of sophisticated cyberattacks, including booting, sinkhole, cloud malware, access control attacks. Based on their study, the authors outline the corresponding future research directions and associated problems within the scope of their study. Xiao et al. [42] examined various cyberattacks against IoT environments, such as eavesdropping, spoofing, and jamming. In addition, the authors identify several IoT-specific security techniques based on learning, such as malware identification, access control and secure offloading. The study discusses the challenges related to state-of-the-art machine learning-based protection techniques, such as computing and communication overhead and partial state observation. Tahsien et al. [43] presented a discussion about the layered IoT architecture. The importance of IoT security in terms of possible attacks under different types, such as physical and cyber attacks, attack surfaces including device perception and cloud applications, and the effects of such attacks including accessibility, integrity, and authorization are discussed in detail. In addition, prospective ML-based contributions to IoT security based on different ML and DL algorithms are presented.
III-A3 Vision For 6G networks
In a prospective overview that outlines the principles of a 6G system, Saad et al. [4] position 6G as a fundamental transformation of Self-Organizing Networks (SONs), whereby the network only adjusts its operations to particular environmental states, into a Self-Sustaining Network (SSN) capable of sustaining its Key Performance Indicators (KPIs) across the extremely complex and highly dynamic operating environments arising from the rich application landscape of 6G. The authors elaborate that AI and specifically Reinforcement Learning (RL) address the goal of building SSNs that can independently sustain high KPIs and handle network resources, functionality, and oversight. Furthermore, in the same study, it is anticipated that the AI-based 6G features to be joined by a collaborative network intelligence located at the edge, resulting in a 6G system that can accommodate future services such as Massive URLLC, and may even be capable of replacing classical network frame structures.
Although the main constituents of the 6G architecture remain undefined and yet to be standardized, some aspects can be foreseen and their associated threats are being discussed. For example, Siriwardhana et al. [5] provided a future-oriented view of the immense role of AI in 6G network security, pinpointing upcoming research directions through its discussion of the AI-based security and privacy challenges, along with some proposed possible solutions. The authors broadly categorize the 6G network threat landscape into two categories, namely, architectural and technological threats. The first entails attacks on the infrastructural level such as attacks on the User Plane Micro Services (UPMS), and the Control Plane Micro Services (CPMS), while the second entails attacks on embedded technologies such as attacks on ML, blockchain, cryptographic protocols, and Visible Light Communication (VLC).
III-B Machine Learning Associated Threats
Although ML-based solutions can provide reasonable defense capabilities, they are prone to various adversarial attacks. Motivated by the facts above, we present different studies on ML threats with respect to three learning paradigms, as presented below:
III-B1 Centralized Edge Learning
Centralized learning is considered the first learning paradigm for ML, where data is gathered at one location for training. Despite its advantages, This approach also involves disadvantages and associated threats. For instance, Liu et al. [45] review the security-related threats to ML and provide a methodical study on such threats in two dimensions: 1) training stage and 2) test/inference stage. Then, the authors classify existing protective techniques used for ML into four classes, including, assessment techniques, training phase precautions, inference-test stage defenses, and privacy and data protection precautions. In another study, Xue et al. [46] present threat models where ML classifiers are targeted by adversaries. The analysis of the reasons behind the possibility of being attacked is presented. Therefore, security concerns are categorized under five categories, including, poisoning of the training set, backdooring through the training set, adversarial attacks, model theft, and inference attacks. In addition, a number of suggestions on ML-related security evaluations are also provided.
Hu et al. [47] present the entire life-cycle of an AI-based operating environment as a roadmap to outline potential security-related threats occurring at each phase, and subsequently elaborate on the corresponding countermeasures that can be taken. Oseni et al. [8] and Liu et al. [48] focus on investigating adversarial attacks against AI-based systems, taking into account areas such as available methods for the generation of adversary samples. Oseni et al. also expands on the mathematical engines of AI, particularly the emerging variants of reinforcement and federated learning in order to illustrate how vulnerabilities in AI models are exploited. In addition, the study considers several cyber defenses to help prevent AI systems against these types of threats, including, data sanitization, robust statistics, defensive distillation, and gradient masking.
Kuzlu et al. [9] provides a study in which concepts around IoT security are presented, with a focus on attacks using and targeting AI. The authors provide a classification of AI attacks into three categories, namely 1) vulnerability scanning automation, which includes fuzzing and symbolic execution, 2) input attacks, and 3) data poisoning and fake data insertion, involving datasets and algorithm poisoning. Hao and Tao [49] review existing adversary evasion and poisoning attacks in smart grids. The types of adversarial examples are classified based on the aspect of the threat, including the attacker’s influence, knowledge, specificity, computation and approach, and security breach. Multiple attack types per category are also provided, such as white box/black box attacks, targeted/non-targeted attacks, and specific approaches used such as gradient, decision, and transfer attacks. The authors propose six approaches for effective checks to mitigate adversarial examples on image classification, consisting of the following: gradient hiding, adjoint detection models, statistical methods, preprocessing methods, the ensemble of classifiers, and proximity metrics.
Chen et al. [50] discussed the critical infrastructure frameworks monitored by the IoT and the associated security vulnerabilities with the main focus on Advanced persistent threat (APT) attack patterns. Specifically, the authors examine a variety of cutting-edge AI-based approaches to discover and successfully mitigate attacks on such networks. Among them, 14 AI-based approaches are selected according to their application frequency. Al-Rubaie et al. [51] focused on different threats to ML privacy such as reconstruction, de-anonymization, and membership inference attacks. For privacy preservation, the authors consider multiple cryptographic techniques as a defense barrier to mitigate these attacks and preserve ML privacy in its different stages (i.e., data preparation, learning, and inference), including DP, homomorphic encryption (HE), garbled circuits, and secure processors.
III-B2 Federated Edge Learning
The introduction of FL provided a privacy-preserving paradigm for ML training. Although FL can ensure privacy preservation to some extent, there are FL-specific threats and the maintained privacy is only offered under certain circumstances. Mothukuri et al. [19] review specific privacy and security issues in federated learning that need to be addressed. The outcome of their investigation suggests that, overall, there are less significant specific privacy threats associated with federated learning than there are security threats. The authors state the following highest priority security threats: communication blockages, poisoning, and backdoor attacks, whereas inference attacks pose one of the greatest threats to FL privacy. Liu et al. [52] reviewed current threats and defense mechanisms in the FL domain across all stages of FL, namely data and behavioral auditing, training, and inference. For each stage, the authors discussed potential threats, related attacks, and available defenses. Lyu et al. [14] review existing privacy threats and attacks in all categories of FL, namely Horizontal Federated Learning (HFL), Vertical Federated Learning (VFL), and Transfer Federated Learning (TFL). The authors also describe the privacy leakage issues in FL by pointing out the fact that these issues can originate from the aggregator or from individual participants. It is important to note that threats can be associated with both insiders and outsiders, as well as with malicious or semi-honest nodes. The attacks on FL are classified into two broad categories: 1) poisoning attacks and 2) inference attacks. Yang et al. [11] present an overview of federated learning, a novel approach addressing challenges such as data privacy and security in AI development. The authors offer definitions, classifications, and potential applications of a secure federated learning framework, further discussing its successful application across different business domains. The authors advocate for a paradigm shift in AI, redirecting the focus from enhancing model performance (the predominant current focus) to exploring compliant data integration methods, thereby aligning with data privacy and security laws. Wahab et al. [15] provide a tutorial on Federated Learning and its related concepts, technologies, and learning approaches. The authors then categorize the literature into high-level challenges addressed by Federated Learning and further divide these into low-level challenges. This three-tier classification allows a deeper understanding of the topic and the methods employed to tackle particular problems. The paper also offers a set of desirable criteria and future research directions for each category of high-level challenges, with the aim to assist the research community in designing innovative and efficient solutions.
Implementing blockchain as a ledger technology can help decentralize FL training without requiring a central server, improving security and scalability. This combination of FL and blockchain has led to the creation of a new paradigm called FLchain. FLchain potentially transforms Mobile-edge computing (MEC) networks into decentralized, secure, and privacy-enhancing systems. Zhu et al. [36] identify key issues in FL that blockchain can address and categorizes existing system models into decoupled, coupled, and overlapped classes based on federated learning and blockchain integration. It compares the benefits and downsides of these models and investigates potential solutions to their limitations. Issa et al. [35] proposes the use of blockchain technology and smart contracts to secure FL in IoT systems. It reviews blockchain-based FL methods and discusses current IoT security issues. The paper also covers IoT data analytics from a security perspective, as well as the challenges and risks of integrating blockchain and FL in IoT. Open research questions are addressed, and a thorough literature review of blockchain-based FL approaches for IoT applications is provided. Nguyen et al. [16] provides an overview of FLchain’s fundamental concepts and explores its opportunities within MEC networks. Several challenges related to FLchain design are also identified, including communication cost, resource allocation, incentive mechanism, security, and privacy protection. Potential applications of FLchain in popular MEC domains such as edge data sharing, edge content caching, and edge crowdsensing are discussed. Therefore, Alazab et al. [17] present a review of various FL models developed to enhance authentication, privacy, trust management, and attack detection. The article also explores real-time use cases that have recently employed FL to preserve data privacy and enhance system performance. Ghimire and Rawat [20] survey the application of FL, a privacy-aware machine learning model, in enhancing the security of IoT systems. They compare centralized learning, on-site distributed learning, and FL, focusing primarily on the security aspects. The discussion also covers performance issues such as accuracy, latency, and resource constraints that might affect the overall functionality of IoT. In addition to evaluating current research efforts, challenges, and trends in the field, the authors consider future developments in this paradigm. The article provides readers with an in-depth understanding of FL’s role in cybersecurity, outlining different security attacks and countermeasures.
Despite the potential of FL, non-IID (non-independent and identically distributed) data present on individual devices participating in the FL process pose significant issues. This statistical heterogeneity can hamper the model’s performance and discourage user participation in FL. Ma et al. [21] discusses the challenges posed by non-IID data in the context of FL. The authors provide a review of the state-of-the-art solutions for non-IID problems, aiming to fill a gap in the literature and facilitate further implementation of FL. Boobalan et al. [23] propose an overview of the combination of FL and IIoT, addressing data privacy and on-device learning motivations, potential usage of machine learning, deep learning, and blockchain techniques for secure IIoT. It also explores the management of large and diverse data sets and discusses applications in industries like automotive, robotics, agriculture, energy, and healthcare. The upcoming deployment of billions of IoT devices, facilitated by faster Internet speeds from 5G/6G, will produce a massive amount of data, including potentially sensitive user information. This surge in data will escalate communication and storage costs and intensify privacy concerns within traditional, centralized cloud-learning systems for IoT platforms. By eliminating the need for data centralization, FL reduces costs and enhances user-level privacy. However, implementing FL in IoT networks is not without its challenges. Zhang et al. [28] delves into the opportunities and hurdles of integrating FL in IoT platforms and its potential to enable a variety of IoT applications. It specifically identifies and explores seven major challenges of using FL in IoT platforms and underlines some recent promising strategies for overcoming them. In a study by Jere et al. [53], the authors point out client-side attacks as the most significant threats to FL platforms since malicious FL clients are capable of tampering with and repositioning the model edges while developing it. In addition, the authors discuss different critical FL attacks, including Generative Adversarial Network (GAN) reconstruction, data poisoning, membership inference, and model inversion attacks. Some of the defense mechanisms reported in the paper include DP, robust aggregation, and outlier detection. The details of Federated Edge Learning are presented in Algorithm 1.
III-B3 Edge learning
This learning paradigm incorporates learning from multiple sources. Chen et al. [54] discussed a range of possible avenues for deploying a variety of distributed learning approaches on real-world wireless edge networks. The outlined approaches include FL, multi-agent RL, Federated Distillation (FD), and distributed inference. The authors also highlight both the benefits as well as potential security and privacy issues these approaches may encounter, such as gradient leakage in FL. In an attempt to answer the question of whether distributed learning is appropriate for wireless communications, Qian et al. [26] survey the state-of-the-art research on distributed learning for wireless communication, as well as the application cases, framework, algorithms, and other suitable alternatives for distributed learning. The research focuses on three layers: 1) physical, 2) medium access control (MAC), and 3) network layer. Furthermore, other emerging areas such as tensor and blockchain technologies are explored by reporting that these systems are prone to security attacks.
Ma et al. [27] provide an information-exchange level security and privacy risk classification of distributed ML, which is organized according to the core phases of an ML workflow, namely preprocessing, learning, knowledge extraction, and result intermediation. The authors investigate and discuss possible risks involved in each level through an overview of current attack techniques, such as model poisoning/inversion attacks, inference attacks, label leakage, and data reconstruction attacks. The details of Distributed Edge Learning are presented in the Algorithm 2. In the case where a given client is involved in multiple FL sessions, it can communicate its real-time global status which is an array of every FL session status. The global status can be formalized as $GS_{C_{ID}}=[{S_{0}:Status},..,{S_{n}:Status}]$, where $C_{ID}$ presents the client ID, and $S_{ID}$ presents the FL session ID. The notations used in Algorithms 1 and 2 are presented in Table VI.
TABLE VI: Notations used in Algorithms 1 and 2
Notation Description
$\eta$
Learning rate
$Epo$
Number of local epochs
$Batch$
Local minibatch size
$K$
Total clients number
$State$
State of the client
$x$
Single sample
$C$
Fraction
$R$
Global rounds
$S_{t}$
A subset of selected clients
$n$
The number of local samples
$f$
Model
$\mathcal{P}$
Pre-processed dataset
$f_{c}(.)$
Loss function
$InitializeModel(.)$
Random model initialization
$ConnectedClients(.)$
Number of active clients required
$RequestUpdate()$
A function used to request the model updates
$ReceiveUpdate()$
A function used to receive the model updates
$AggregateModels()$
A function used to aggregate the model updates
$UpdateModel()$
A function used to updates the model after the aggregation
Data: $\eta$ , $Epo$, $Batch$, $K$, $k$.
Edge Server EdgeFedLearn ($K$, $C$, $R$):
       $f_{1}\leftarrow InitializeModel()$
       for $t=1,..,R$ do
             $S_{t}$leftarrow Subset(max($C\cdot K,1$), "$random"$)
             Parallel.for $k$in $S_{t}$  do
                   $f_{t+1}^{k}$leftarrow $ClientUpdate(f_{t},k$)
             end for
            $f_{t+1}$leftarrow $\sum_{k=1}^{K}\frac{n_{k}}{n}f_{t+1}^{k}$
       end for
      Broadcast $f_{t+1}$ the updated model to clients
IoT device  ClientUpdate ($f$, $k$):
      1 $\mathcal{B}$leftarrow Split($\mathcal{P}$, $Batch$)
      2 for i = 1,..,$Epo$ do
             for $b$in $\mathcal{B}$ do
                   $f$leftarrow $f$- $\eta\nabla f_{c}(x,b)$
             end for
       end for
      Return $f$ to Edge Server
Algorithm 1 Federated Edge Learning
Data: $\eta$, $Epo$, $Batch$, $K$ , $k$, $State$
$f_{1}\leftarrow InitializeModel()$ ;
Broadcast $f_{1}$ the initial model to clients
while $K$neq length(ConnectedClients()) do
       continue
end while
$RequestUpdate()$
$ReceiveUpdate()$
$AggregateModels()$triangleright State =1
$UpdateModel()$triangleright State =2
$State=1$ EdgeFedLearn ($K$, $C$, $R$):
       for $t=1,..,R$ do
             $S_{t}$leftarrow Subset(max($C\cdot K,1$), "$random"$);
             Parallel.for $k$in $S_{t}$  do
                   $f_{t+1}^{k}$leftarrow $ClientUpdate(f_{t},k$)
             end for
            $f_{t+1}$leftarrow $\sum_{k=1}^{K}\frac{n_{k}}{n}f_{t+1}^{k}$
       end for
      Broadcast $f_{t+1}$ the updated model to clients
$State=2$ ClientUpdate ($f$, $k$):
      1 $\mathcal{B}$leftarrow Split($\mathcal{P}$, $Batch$) ;
      2 for i = 1,..,$Epo$ do
             for $b$in $\mathcal{B}$ do
                   $f$leftarrow $f$- $\eta\nabla f_{c}(x,b)$
             end for
       end for
      Return $f$ to Client
Algorithm 2 Distributed Edge Learning
TABLE VII: Difference between Edge Learning in 4G IoT, Edge Learning in 5G IoT, and Edge Learning in 6G IoT.
Bandwidth
10 Mbps
1 Gbps
10 Gbps
Latency
100 ms
1 ms
100 µs
Data Rate
100 MBps
10 GBps
100 GBps
Energy Efficiency
Medium
High
Very High
Network Density
1000-10,000 devices per square km
100,000-1,000,000 devices per square km
10 million+ devices per square km
Network Connectivity
Cellular network and Wi-Fi hotspots
Integrated satellite, airborne, and terrestrial networks
Spectrum Efficiency
2-3 bps/Hz
10-20 bps/Hz
100 bps/Hz
Network Architecture
Centralized
Decentralized
Autonomous
Applications
Streaming, Video Conferencing, Gaming
IoT, Autonomous Vehicles, Smart Cities
Interoperability
Limited
Improved
Full Interoperability
Edge Computing Capability
Limited (Basic analytics, traditional ML)
Moderate (Advanced analytics, deep learning)
Advanced (Distributed computing, edge AI)
AI Integration
Limited
Advanced
Full AI Integration
Virtual Reality Integration
Limited
Advanced
Full Virtual Reality Integration
Autonomous Devices
Limited
Advanced
Full Autonomous
Coverage
Limited (Urban areas)
Wide (Urban and rural areas)
Very wide
Reliability
Medium (Packet loss, interference)
High (Low packet loss, low interference)
Ultra-high (No packet loss, no interference)
Semantic Communications
Holographic MIMO Surfaces
Not supported
Limited support for holographic MIMO
Millimeter-Wave and Terahertz Bands
Not supported or limited support
Increased support for millimeter-wave and terahertz bands
Network Slicing
Not supported
Physical Layer
Orthogonal frequency division multiplexing (OFDM)
Massive MIMO
Not Available
Available
Available
uMUB
Not Available
Available
Available
uHSLLC
Not Available
Available
Available
Near Space Communications
Not supported
Not supported
Full integration for near space communication and data transmission
mMTC
Available
Available
Available
uHDD
Not Available
Not Available
Available
Key Technologies
LTE, WiMAX, VoLTE, MIMO
Use Cases
Smart homes, wearables, fleet management
Smart factories, autonomous vehicles, e-health
Smart cities, augmented reality, digital twins
III-C Quantitative Comparison of Edge Learning in 4G, 5G, and 6G-enabled IoT Environments
Edge learning is a subset of edge computing, which aims to perform machine learning algorithms on data generated at the network’s edge, i.e., closer to the source of data generation. Figure 5 illustrate the network architecture and communication processes for centralized edge learning-6G-enabled IoT and federated edge learning-6G-enabled IoT. Centralized edge learning and federated edge learning are two approaches to training machine learning models in 6G-enabled IoT, which differ in how they handle data. In centralized edge learning, all the data used to train the model is collected and stored in a central location, such as an edge server or cloud-based platform. The model is then trained based on this centralized dataset, which may be relatively large and diversified. On the other hand, federated edge learning refers to training machine learning models on distributed datasets across many devices. This approach can be particularly useful when data privacy is a concern, as the data remains on the device and is not transmitted to an edge server. However, both approaches have advantages and disadvantages, and the suitable approach will depend on the specific needs and constraints of the current machine learning task for 6G-enabled IoTs [55, 56].
With the advent of IoT devices, there is an ever-increasing demand for real-time data processing and analysis, which is fueling the need for edge learning [57]. Table VII compares the features of Edge Learning in 4G IoT, 5G IoT, and 6G IoT. We can observe a significant improvement in various aspects of each generation of technology. Moving from 4G to 6G, we observe significant improvements in latency, bandwidth, the number of devices supported, computational power, energy consumption, and cost per device. These improvements are expected to profoundly impact the applications of edge learning [33, 30, 29]. For instance, the bandwidth, data rate, energy efficiency, and network density in 6G IoT are significantly higher compared to 4G IoT. Moreover, the latency in 6G IoT has improved to 100 microseconds compared to 100 milliseconds in 4G IoT. The level of AI and Virtual Reality integration in 6G IoT is also more advanced compared to 4G IoT. Furthermore, the level of autonomous devices has improved greatly in 6G IoT. Despite this, the availability of Massive MIMO remains constant throughout the three generations of the technology. However, other technologies like uMUB and uHDD are only available in 5G IoT and 6G IoT. The level of interoperability has also improved greatly in 5G IoT and 6G IoT.
III-D Practical Deployments of AI in Edge/Fog
Although edge computing and fog computing are frequently used as synonyms, they are not identical concepts. Both emphasize the distribution of computing resources near the data’s origin and usage; however, their methods differ. Edge computing involves processing data at or near the point of generation, like IoT devices, sensors, or other endpoints. In contrast, fog computing is a broader notion that expands the cloud computing paradigm to the network’s periphery. This involves distributing computing, storage, and networking resources across multiple layers, from the edge devices to the cloud [62]. Effectively managing IT assets in cloud and fog/edge environments is a challenging task that necessitates a methodical decision-making process. The scarcity and heterogeneity of resources, in conjunction with the dynamic and diverse workloads, along with the unpredictable nature of advanced IT environments, have compounded the complexities of managing resources in such a landscape. It is crucial to observe the infrastructure’s behavior since comprehending the workload’s behavior can mitigate the intricacy of the challenge and improve the outcome of a particular implementation [63]. To overcome these challenges, adopting AI/ML-based solutions has gained traction, leveraging their ability to make sequential decisions and achieve optimal outcomes in this complex and ever-changing environment [64]. For example, in anticipating workload patterns or spatiotemporal impacts ahead of time for assisting in orchestrating resources [65].
AI-enabled fog and edge computing are being deployed in various industries, including healthcare, transportation, manufacturing, and retail, among others. Edge AI technology has been developed and commercialized by various companies, and there is a wide range of industrial products that are currently available or in development. Such implementations include edge AI chips, where companies such as NVIDIA [66], Intel, and Qualcomm have developed specialized chips designed to run AI algorithms on edge devices. These chips are used in a variety of applications, including autonomous vehicles, drones, and smart cameras. Also, Edge AI software platforms, where several companies, such as AWS IoT Greengrass, Google Cloud IoT Edge [67], and Microsoft Azure IoT Edge, have developed software platforms that enable developers to deploy AI models on edge devices. These platforms provide tools for building, training, and deploying machine learning models. In addition, edge AI cameras with built-in AI capabilities are becoming increasingly popular in industrial and commercial settings [68]. These cameras are used for applications such as facial recognition, object detection, and anomaly detection. Also, edge AI sensors, where IoT sensors with built-in AI capabilities are also becoming more common. These sensors are used in applications such as predictive maintenance, asset tracking, and environmental monitoring. Furthermore, Edge AI robots, where robots with onboard AI capabilities are being used in a variety of industrial applications, including warehouse automation, manufacturing, and agriculture. In terms of business models, AI-enabled fog and edge computing are being deployed through various models, including software as a service (SaaS), platform as a service (PaaS), and infrastructure as a service (IaaS). For instance, in the healthcare industry, companies such as Philips offer edge computing as a service for remote patient monitoring [69]. Similarly, in the manufacturing industry, companies such as GE are offering predictive maintenance as a service through their Predix platform [70].
On the other hand, AI needs a fair amount of computing power and appropriate energy backing for learning. However, devices often fail to meet both of these requirements. One solution is to do all the heavy lifting in the cloud, which poses new challenges, including a shortage of bandwidth and elevated latency when dealing with a wide range of different AI models running on a wide range of endpoint devices. The emergence of edge/fog computing enables AI to be distributed close to end devices and users at the edge and on the endpoint with some processing and storage capabilities, addressing the demands for both high network steadiness and minimal latency. In another study conducted by Joshi et al. [71], an emerging concept of edge intelligence, highlighted by the authors, was the so-called "all-in-edge" tier, in which the formation and inference of AI models are carried out only by the edge servers. Research on AI-driven optimization of edge/fog systems can be roughly classified into two major categories: optimization-related challenges (e.g., offloading, resource allocation [72], energy consumption [73]) and privacy/security-related challenges [32]. For instance, DRL has been used to effectively learn the network dynamics in the work of. Cheng et al. [74]. The authors proposed a DRL-based offloading method based on an embedded space-air-ground network to minimize energy and latency overhead.
[…]

### v2

@@ II-D Quality assessment
In our study, we conducted a quality assessment using a questionnaire that included ten questions related to the research questions and objectives. These questions covered a range of topics, including 1) the clarity and appropriateness of the research questions, 2) the comprehensiveness of the literature review, 3) the appropriateness and relevance of the datasets used, 4) the effectiveness of the classification methods used, 5) the depth and comprehensiveness of the discussions on proposed solutions, 6) the degree to which the proposed solutions are supported by experimental analysis, and 7) the clarity of the presentation of the results. By answering these questions, we could evaluate the quality and identify the advantages and disadvantages of each work. The formulation of the quality assessment criteria in Table III is based on eight questions that assess the quality of our research concerning the research questions.
III 6G–IoT Intelligence: an overview
For future 6G and beyond, there will be a heavy reliance on Network Intelligence (NI) with AI models representing its backbone [55], which ensures that network management will be adequately managed and fully automated. The AI models have shown considerable effectiveness in solving complex tasks, such as deriving complicated associations from massive, overlapping data, which will be even more required for future sophisticated networks and applications, such as IoE [3]. In this section, we present a vision of the expected effectiveness of intelligence in 6G-IoT systems from three angles as illustrated in Figure 5: 1) the architectural landscape, 2) technological prospects, and 3) security and privacy benefits.
III-A Architectural Landscape
Using AI, specifically, DL provides an innovative path to engineer and improve 6G networks throughout the physical and core layers. Advances in 6G wireless theory and communication will also influence the advancement and expansion of AI in the form of new learning theories and architectures, creating a positive feedback loop, and transforming the wireless landscape, from connecting objects to connecting intelligence [55].
III-A1 Physical (PHY) Layer Enhancements
During the last few years, investigations and efforts have been made to deploy AI in the physical layer of wireless communication networks, including User Equipment (UE) and at the cell edge [56, 57, 55]. Therefore, several problems with current communication systems remain unaddressed due to inaccurate models or non-linearity [56], such as reciprocity in frequency division duplexing (FDD), predicting channels, detecting and reducing interference. According to Ali et al. [56], many of the physical layer’s optimization problems, such as spectrum sensing, optimal beamforming formulation, and throughput maximization employing power control, are non-convex. Such issues may be addressed through dual decomposition techniques that require time-consuming iterative algorithms. On the other hand, DL techniques have significant capabilities in managing such challenges in real-time without sacrificing performance. In addition, AI offers a new way to design the 6G radio interface by further improving the radio environment and communication algorithms. Letaief et al. [55] proposed innovative solutions such as joint source-channel coding (JSCC), task-oriented communication, and semantic communication.
A Reconfigurable Intelligent Surface (RIS) is a customizable surface design capable of manipulating electromagnetic (EM) wave reflections by adjusting electric and magnetic characteristics. The core idea behind RIS is to control the electromagnetic properties of the environment. This manipulation is inherently a physical phenomenon, which makes RIS a tool for the PHY layer. Chen et al. [23] delves into the emerging significance of reconfigurable intelligent surfaces in the context of 6G wireless communication systems, specifically their potential in data delivery and accurate positioning for IoT. RIS can adeptly control radio waves, so its role in 6G communication is growing. This article focuses on the burgeoning interest in RIS-assisted positioning, detailing its working principle, channel model, and characteristics.
III-A2 Medium Access Control (MAC)
ML supports an automated learning paradigm shift towards high-performance algorithms for addressing resource allocation challenges within wireless networks [57, 3]. MAC layer-related tasks such as selecting and matching users for multiple-input multiple-output (MIMO), modulation, and handover control can also be optimized using AI [3, 58]. For instance, given that RL can handle combinatorial action spaces with multi-agent settings, RL is considered suitable for missions where the network can adapt to varying conditions while learning ideal strategies [56]. Therefore, AI plays a vital role in 6G MAC layer optimization tasks for the following ML-based predictive tasks: 1) optimal resource allocation (e.g., in non-orthogonal multi-access (NOMA) and massive MIMO (mMIMO)), 2) ML-based predictive power management (e.g., traffic forecasting and prioritized packets separation), 3) FL-based mobility prediction (e.g., federated echo state network), 4) mobile data offloading, 5) link adaptation, and 6) caching [3, 56, 59, 60].
III-A3 Core-Network and Services Intelligence
One of the key benefits of using AI at the core and edge of the network, specifically by integrating training capabilities into the network nodes, is to enable intelligent data collection, processing, delivery, and utilization at the network edge. These benefits will improve the network service quality. Therefore, the inference method can produce reliable and cost-effective services [59]. As one of several cases of this type, device-server co-inferencing can overcome existing traffic and processing constraints by spreading a sizeable DNN network across different edge node types (I.e., devices and servers) [55]. In addition, future innovative self-organizing and self-repairing characteristics, autonomous connections between devices, and the human-centric network with enhanced intelligence within the 6G network translate directly into a wealth of other communication services [3].
III-A4 6G architecture for energy-efficient communication
Efficient energy use in communication has garnered interest from various sectors, including industrial automation, healthcare, transportation, and more. Furthermore, the surge in AI-driven sixth-generation (6G) technologies aimed at advancing smart automation systems is now in the spotlight for scholars and businesses. The backbone of most advanced automation systems comprises IoT-based user terminal (UT) devices that deliver multimedia content — such as video, audio, image, and text — with precision and efficacy. Sodhro et al. [17] address the growing importance of energy-efficient communication in various sectors and the rise of AI-driven 6G technology in intelligent automation systems. These systems predominantly use IoT devices for high-quality multimedia content delivery. Recognizing that the traditional Quality of Service (QoS) doesn’t adequately reflect user sentiment during multimedia transmission, the study focuses on Quality of Experience (QoE). The paper introduces a QoS-based joint energy and entropy optimization algorithm (QJEEO), formulates a 6G-influenced multimedia data structure for QoE evaluation, establishes a link between surveyed user scores and IoT device performance metrics, and devises a correlation model for QoS and estimated QoE perceptions. Experimental findings underscore that QoE can be assessed and aligned with QoS factors, specifically packet loss ratio and transfer delay, to enhance user satisfaction in energy-efficient, 6G-based multimedia transmissions.
III-A5 5G/6G IoT testbeds
The table XI provides a list of 5G and 6G IoT testbeds that are crucial for testing and developing advanced 5G and 6G IoT technologies and applications. The testbeds are designed to test and develop 5G/6G-based IoT services and applications, such as smart transportation, energy management, healthcare, and industrial automation. The European Union has funded several testbeds listed, including 5G-TRANSFORMER, 5G-MoNArch, 5G-PICTURE, 5G-CORAL, and 5G-EmPOWER. Other funded testbeds include the 5G Open Innovation Lab in the USA, ENCQOR 5G in Canada, 5G City in Denmark, the 5G Alliance for Connected Industries and Automation in Germany, and the 6G platform in Germany. The testbeds feature network slicing, multi-access edge computing (MEC), cloud computing, virtualization, terahertz and sub-terahertz communications, and artificial intelligence (AI).
III-B Supported Technological Prospects
[…]